OpenStudy (anonymous):

Find the linearization of the function f(x,y) = \sqrt{100 - 4 x^2 - 2 y^2} at the point (-4, 4). L(x,y) = Use the linear approximation to estimate the value of f(-4.1, 4.1) =

OpenStudy (anonymous):

\[ \Delta L = m_x\Delta x + m_y\Delta y \]

OpenStudy (anonymous):

I believe we start with this.

OpenStudy (anonymous):

Then we can expand it a bit...

OpenStudy (anonymous):

\[ L(x,y) - L(x_0,y_0) = m_x(x-x_0)+m_y(y-y_0) \]

OpenStudy (anonymous):

Now \(\langle m_x,m_y\rangle\) is just the gradient.

OpenStudy (anonymous):

Basically \(\langle m_x,m_y\rangle = \langle f_x(x_0,y_0), f_y(x_0,y_0)\rangle \)

OpenStudy (anonymous):

\[ L(x,y) - L(x_0,y_0) = f_x(x_0,y_0)\cdot (x-x_0)+f_y(x_0,y_0)\cdot (y-y_0) \]

OpenStudy (anonymous):

So, find \(f_x\) and \(f_y\). Remember that in this case \((x_0,y_0) = (-4,4)\) and \((x,y) = (-4.1,4.1)\). That accounts for everything.

OpenStudy (anonymous):

Remember that the single variable version is: \[ \Delta y \approx dy = y'dx \]In this case we are using: \[ \Delta z \approx dz = z_xdx+z_ydy \]